OpenStudy (anonymous):

By what factor would the mass of a spaceship increase if it accelerates from rest to 0.750c? If the rest mass of the spaceship is 25.0 × 103 kg, then calculate its kinetic energy when travelling at 0.750c according to: a) Special Relativity and b) Classical Physics

OpenStudy (anonymous):

\[m(moving) = \frac{m(rest)}{\sqrt{1- \frac{v^2}{c^2}}}\]

OpenStudy (vincent-lyon.fr):

@Mashy Just to illustrate "relativity" of concepts, "moving mass" does not exist in French textbooks or teaching! Mass is rest-mass, full stop: Mass does not change when a body accelerates. Different perspective on same theory. Funny, isn't it?

OpenStudy (anonymous):

well.. i m pretty sure they don't consider moving mass unless they start talking about einstiens mass energy relationship and relativity and sub luminal speeds.. obviously when they talk about elementary mechanics they would neglect the relativistic mass and stuff :).. else people won't understand anything.. !

OpenStudy (anonymous):

Well if you could literally accelerate a rocket from rest to 0.750c from earth into, you can be the next Stephen Hawking...

OpenStudy (anonymous):

no you would be shish kebab :P

OpenStudy (anonymous):

or me ;-)

OpenStudy (vincent-lyon.fr):

@Mashy No, I'm talking 100% relativity. We have no such thing as mass increase with speed. But we have relativistic definition of momentum. So in the end, the theory is the same.

OpenStudy (anonymous):

wow.. like that.. :O.. i see.. i m not really much into relativity.. so i don't really know how its taught here anyways.. but when you talking about quantum mechanics.. especially sommerfields model how he improved the Bohr's model.. he considered the relativistic variation of the mass of the electrons.. so how did they explain that part? again mentioning it as 'momentum' ?